# The triton lifetime from nuclear lattice effective field theory

### v1

@@ 1 Introduction
In case of the two-pion exchange, these LECs that are called $c_{1,3,4}$. They can be precisely determined from pion-nucleon scattering [8], demonstrating the power of chiral symmetry in connecting seemingly unrelated processes. To pin down the LECs $c_{D}$ and $c_{E}$, that parameterize the one-pion exchange and contact term topologies at this order, respectively, one must consider observables in three-nucleon systems. In a first systematic analysis of neutron-deuteron scattering at N2LO, these LECs were fixed from the triton binding energy and the neutron-doublet scattering length [9], although these quantities display some correlation. It was pointed out first in [10] that triton $\beta$-decay, that is the triton lifetime, together with the binding energies in the $A=3$ system can lead to a robust determination of $c_{D}$ and $c_{E}$. It could later be shown that using the cross section minimum in low-energy proton-deuteron as well as neutron-deuteron scattering can also leads to a fairly precise determination of the short-distance LEC $c_{D}$, see [11] (and references therein). It is worth pointing out that in these continuum approaches, a completely consistent regularization scheme for two- and three-body forces as well as external currents has only become available very recently [12]. In pionless EFT, where all interactions are represented by contact terms, triton $\beta$-decay has been studied in [13].
@@
Here, we approach the problem of triton $\beta$-decay from a different perspective, namely in the framework of Nuclear Lattice Effective Field Theory (NLEFT). For an introduction to that method, see e.g. [14, 15]. This lattice approach has proven successful in solving problems that were previously considered intractable with conventional methods, like the first ab initio calculations of the Hoyle state in the spectrum of 12[16] and of alpha-alpha scattering [17]. These calculations were performed at N2LO on a coarse lattice, which limits the theoretical precision. A major step forward in achieving high precision in NLEFT for many-nucleon systems was recently made using the so-called the wavefunction matching method at next-to-next-to-next-to-leading order (N3LO) in the chiral expansion [18]. High-fidelity chiral interactions at N3LO often encounter significant sign problems due to the cancellation of positive and negative contributions, making Monte Carlo calculations impractical. The wavefunction matching method introduced in Ref. [18] resolves this issue for two-nucleon interactions at N3LO. This method was successfully applied to light nuclei, medium-mass nuclei ($A\leq 58$), neutron matter, and nuclear matter, and good agreement with the empirical data was found. Despite the success of the wavefunction matching method in improving theoretical precision, the method’s application thus far has been restricted to first-order perturbation theory. This limitation means that triton $\beta$-decay calculations at N3LO, requiring higher-order corrections for realistic wave functions, cannot be directly performed using the methods from Ref.[18]. One potential solution to this problem is to extend calculations to second-order perturbation theory. Recent advances in perturbative quantum Monte Carlo (QMC) calculations, as detailed in Ref. [19], offer a viable approach for computing these perturbative corrections. Alternatively, fully non-perturbative calculations for three-nucleon systems can be performed to obtain realistic wave functions at N3LO for triton $\beta$-decay calculations. In this paper, we employ the latter approach.
@@
In our calculations, we use the same action as in Ref. [18]. Our findings indicate that for a given set of LECs, the theoretical precision aligns well with experimental data, allowing us to explore variations of the smeared $c_{D}$ and $c_{E}$ LECs and their impact on the triton lifetime. It should also be stressed that the triton $\beta$-decay offers a benchmark for investigating the weak interactions in nuclei, eventually paving the way for the calculation of the neutrinoless double-$\beta$ decay in nuclei like e.g. 48Ca or 76Ge. Employing wavefunction matching and perturbative quantum Monte Carlo methods, such nuclei can now be tackled within NLEFT.
@@ 2 Formalism
$\langle GT\rangle_{\text{emp}}=\sqrt{\frac{(K/G_{V}^{2})/[(1+\delta_{R})\,t_{1/2}\,f_{V}]-\langle F\rangle^{2}}{({f_{A}}/{f_{V}})\,g_{A}^{2}}}=1.6497\pm 0.0023\,.$ (5)
@@
Our approach to triton $\beta$-decay calculations employs NLEFT, and we briefly discuss the main ingredient of the method used here and refer to Ref. [18] for more details. In a nutshell, a new quantum many-body approach, the so-called wave function matching, transforms the high-fidelity interaction between particles so that the wave functions of the high-fidelity Hamiltonian up to some finite range match that of an easily computable Hamiltonian. More precisely, wavefunction matching operates entirely in the two-nucleon sector. For the nuclear case, this simplified Hamiltonian consists of Wigner SU(4) symmetric two-nucleon forces and the properly regularized one-pion exchange, and it is treated fully non-pertubatively. To bring the chiral Hamiltonian $H$ close to the simplified Hamiltonian $H_{S}$, a unitary transformation is performed leading to $H^{\prime}=U^{\dagger}HU$, and the differences to the full chiral Hamiltonian, $H^{\prime}-H_{S}$, are then calculated in first order perturbation theory. Finally, fitting the various locally and non-locally smeared 3NF operators to the nuclear binding energies with $3\leq A\leq 58$, one can predict the corresponding nuclear charge radii as well as the equation of state of pure neutron as well as nuclear matter. All of these quantities agree with the data. In this work, we extend beyond the first-order perturbation theory previously used for calculations at N3LO in the chiral effective field theory [18], applying fully non-perturbative calculations to three-nucleon systems, with the corresponding Hamiltonian given in A. This approach allows us to obtain realistic wave functions necessary for accurate triton $\beta$-decay calculations at N3LO. Our method involves solving the Schrödinger equation for the three-nucleon system using the Lanczos eigenvector method [28]. The three-dimensional space is represented by a finite volume $L\times L\times L$, with $L$ the spatial extension, and with lattice spacing $a=1.32$ fm in the spatial direction. This corresponds to the magic momentum cutoff of $p_{\rm max}\simeq 465\,$MeV that best displays the hidden spin-isospin symmetry of QCD [29].
@@ 3 Results and discussion
As noted before, we employ the same lattice action as in Ref. [18], with the key difference being the non-perturbative treatment of the interactions. Our calculations start with the determination of the ground state energies of 3H and 3He. Using the NLEFT framework, we obtain infinite-volume extrapolated ground state energies of $8.33(2)$ MeV and $7.62(2)$ MeV, respectively, as shown in the left panel of Fig. 2. These values are slightly below the empirical values of $8.48$ MeV and $7.72$ MeV, respectively. Since the energies at $L=12$ (15.84 fm) are very close to the extrapolated results, we present the results at $L=12$ for all further analyses.
Figure 2: The ground state energies of 3He and 3H nuclei as a function of the box size $L$ (left panel). The trajectory of $f_{D}$$f_{E}$ that leaves these energies invariant (right panel). The dark (light) red/blue bands refer to the 90 (99) % confidence level of the extrapolation. In the right panel the red squares refer to the actual calculated combinations and the blue line is a 3rd order spline interpolation.
@@
where the explicit forms of $V_{c_{D}}^{(n)}$ and $V_{c_{E}}^{(n)}$ are given in Appendix A. It is important to note that all these potential terms, with their various forms of smearing, are associated with the LECs $c_{D}^{(n)}$ and $c_{E}^{(n)}$, also listed in Appendix A, making a direct comparison with the continuum values of these LECs impossible. Here, in Eqs. (6) and (7), we introduce the factors $f_{D}$ and $f_{E}$ to allow us to vary the overall contributions from the one-pion-exchange potential $V_{c_{D}}$ and the SU(4) symmetric short-range potential $V_{c_{E}}$. We then explore the trajectory of $f_{D}$ versus $f_{E}$ that maintains the binding energies of the three-body systems invariant [9, 10].
In the following analysis, we vary $f_{E}$ within the range $[0.4,1.4]$, which, due to the aforementioned correlation, causes $f_{D}$ to vary from $0.7$ to $1.0$, as shown in the right panel of Fig. 2. Because the potential terms in Eqs. (6) and (7) respond differently to the three-body systems, the trajectory observed in Fig. 2 is more complex than the one presented in Ref. [10].
In this study, the infinite-volume extrapolation of the Fermi and Gamow-Teller matrix elements is not considered, and the Fermi operator as well as the Gamow-Teller operator are simultaneously obtained within the consistent framework of NLEFT, as shown as a function of $f_{E}$ in Fig. 3.
@@
Figure 3: Plots of Fermi (left panel) and Gamow-Teller (right panel) matrix elements at $L=12$ (l.u.) as a function of the parameter $f_{E}$ as defined in Eq. (7). The red squares are the calculated values. The grey band in the right panel refers to using the averaged empirical values together with the calculated value of $F$ (left panel). The blue lines are 3rd order spline interpolations.
The final results at $L=12$ are,
$\langle F\rangle$ $=$ $0.999571(48)\,,$ (8)
$\langle GT\rangle$ $=$ $1.6770(34)\,.$ (9)
@@
The uncertainty estimate is based on the variation of the operators over the range of $f_{E}$ and $f_{D}$, as shown in Fig. 3, encompassing sets of three-body forces that describe the three-body systems equally well. A similar dependence on the three-body forces was observed in Ref. [10]. Notably, our value of $\langle F\rangle$ is close to the one given in Eq. (3). The corresponding lifetime is given by
$(1+\delta_{R})\,t_{1/2}\,f_{V}=1102.07(19)~{}\text{s},$ (10)
@@
where the uncertainties are inherited from $K/G^{2}_{V}$. Additionally, we calculate $\langle GT_{F}\rangle$ using $\langle F\rangle$ from Eq. (8) and the averaged empirical half-lives from Refs. [23, 24], injected into Eq. 2, giving (see also the right panel in Fig. 3),
@@ Appendix A Hamiltonian at N3LO
with $\Lambda_{\pi}=300$ MeV the regulator parameter and $M_{\pi}=134.98$ MeV the pion mass. Also, $V_{\rm Coulomb}$ is the Coulomb interaction, $V_{\rm 3N}^{\rm Q^{3}}$ is the 3N potential, $V_{\rm 2N}^{\rm Q^{4}}$ is the 2N short-range interaction at N3LO, $W_{\rm 2N}^{\rm Q^{4}}$ is the 2N Galilean invariance restoration (GIR) interaction at N3LO. For the details of the Coulomb interaction and the two-nucleon (2N) short-range interactions we refer the reader to Ref. [31].
The three-nucleon (3N) interactions at ${\rm Q^{3}}$ consist of locally smeared contact interactions, one-pion exchange interaction with that the two-nucleon contact terms are smeared locally, two-pion exchange potential [7, 9, 34], and two additional SU(4) symmetric potentials denoted by $V_{c_{E}}^{(l)}$ and $V_{c_{E}}^{(t)}$. Therefore, the three-nucleon interactions at ${\rm Q^{3}}$ has the form

### v2

@@ 1 Introduction
In case of the two-pion exchange, these LECs that are called $c_{1,3,4}$. They can be precisely determined from pion-nucleon scattering [8], demonstrating the power of chiral symmetry in connecting seemingly unrelated processes. To pin down the LECs $c_{D}$ and $c_{E}$, that parameterize the one-pion exchange and contact term topologies at this order, respectively, one must consider observables in three-nucleon systems. In a first systematic analysis of neutron-deuteron scattering at N2LO, these LECs were fixed from the triton binding energy and the neutron-doublet scattering length [9], although these quantities display some correlation. It was pointed out first in Ref. [10] that triton $\beta$-decay, that is the triton lifetime, together with the binding energies in the $A=3$ system can lead to a robust determination of $c_{D}$ and $c_{E}$. It could later be shown that using the cross section minimum in low-energy proton-deuteron as well as neutron-deuteron scattering can also leads to a fairly precise determination of the short-distance LEC $c_{D}$, see Ref. [11] (and references therein). It is worth pointing out that in these continuum approaches, a completely consistent regularization scheme for two- and three-body forces as well as external currents has only become available very recently [12]. In pionless EFT, where all interactions are represented by contact terms, triton $\beta$-decay has been studied in Ref. [13].
@@
Here, we approach the problem of triton $\beta$-decay from a different perspective, namely in the framework of Nuclear Lattice Effective Field Theory (NLEFT). For an introduction to that method, see Refs. [14, 15]. This lattice approach has proven successful in solving problems that were previously considered intractable with conventional methods, like the first ab initio calculations of the Hoyle state in the spectrum of 12[16] and of alpha-alpha scattering [17]. These calculations were performed at N2LO on a coarse lattice, which limits the theoretical precision. A major step forward in achieving high precision in NLEFT for many-nucleon systems was recently made using the so-called the wavefunction matching method at next-to-next-to-next-to-leading order (N3LO) in the chiral expansion [18]. High-fidelity chiral interactions at N3LO often encounter significant sign problems due to the cancellation of positive and negative contributions, making Monte Carlo calculations impractical. The wavefunction matching method introduced in Ref. [18] resolves this issue for two-nucleon interactions at N3LO. This method was successfully applied to light nuclei, medium-mass nuclei ($A\leq 58$), neutron matter, and nuclear matter, and good agreement with the empirical data was found. Despite the success of the wavefunction matching method in improving theoretical precision, the calculations in Ref. [18] were carried out using first-order perturbation theory. Since first-order perturbation theory only provides corrections to the energy and not to the wavefunctions, triton $\beta$-decay calculations at N3LO, requiring higher-order perturbative corrections for realistic wave functions, cannot be directly performed using the methods from Ref. [18]. One potential solution to this challenge is to extend the calculations to second-order perturbation theory. Recent advances in perturbative quantum Monte Carlo (QMC) methods, as detailed in Ref. [19], provide an effective framework for incorporating higher-order perturbative corrections, making it particularly well-suited for applications to heavier nuclei. Alternatively, fully non-perturbative methods can be applied to light nuclear systems to generate realistic wave functions at N3LO, as required for triton $\beta$-decay calculations. In this paper, we adopt this non-perturbative approach.
@@
In our calculations, we use the same action as in Ref. [18]. Our findings indicate that for a given set of LECs, our theoretical precision aligns well with earlier theoretical calculations in continuum, allowing us to explore variations of the smeared $c_{D}$ and $c_{E}$ LECs and their impact on the triton lifetime. It should also be stressed that the triton $\beta$-decay offers a benchmark for investigating the weak interactions in nuclei, eventually paving the way for the calculation of the neutrinoless double-$\beta$ decay in nuclei like e.g. 48Ca or 76Ge. By incorporating the wavefunction matching method and advances in perturbative quantum Monte Carlo, NLEFT is now equipped to address these more complex systems, paving the way for high-precision studies of heavier nuclei.
@@ 2 Formalism
$\langle GT\rangle_{\text{emp}}=\sqrt{\frac{(K/G_{V}^{2})/[(1+\delta_{R})\,t_{1/2}\,f_{V}]-\langle F\rangle^{2}}{({f_{A}}/{f_{V}})\,g_{A}^{2}}}=1.6497(23)\,.$ (5)
@@
Our approach to triton $\beta$-decay calculations employs NLEFT, and we briefly discuss the main ingredient of the method used here and refer to Ref. [18] for more details. In a nutshell, a new quantum many-body approach, the so-called wave function matching, transforms the high-fidelity interaction between particles so that the wave functions of the high-fidelity Hamiltonian up to some finite range match that of an easily computable Hamiltonian. More precisely, wavefunction matching operates entirely in the two-nucleon sector. For the nuclear case, this simplified Hamiltonian consists of Wigner SU(4) symmetric two-nucleon forces and the properly regularized one-pion exchange, and it is treated fully non-pertubatively. To bring the chiral Hamiltonian $H$ close to the simplified Hamiltonian $H_{S}$, a unitary transformation is performed leading to $H^{\prime}=U^{\dagger}HU$, and the differences to the full chiral Hamiltonian, $H^{\prime}-H_{S}$, are then calculated in first order perturbation theory. Finally, fitting the various locally and non-locally smeared 3NF operators to the nuclear binding energies with $3\leq A\leq 58$, one can predict the corresponding nuclear charge radii as well as the equation of state of pure neutron as well as nuclear matter. All of these quantities agree with the data. In this work, we extend beyond the first-order perturbation theory previously used for calculations at N3LO in the chiral effective field theory [18], applying fully non-perturbative calculations to three-nucleon systems, with the corresponding Hamiltonian given in A. This approach allows us to obtain realistic wave functions necessary for accurate triton $\beta$-decay calculations at N3LO. Our method involves solving the Schrödinger equation for the three-nucleon system using the Lanczos eigenvector method [28] , which has been successfully applied in nuclear lattice calculations. For a detailed discussion of the method, we refer the reader to Ref. [29], and for its early application in NLEFT, see Ref. [30]. The three-dimensional space is represented by a finite volume $L\times L\times L$, with $L$ the spatial extension, and with lattice spacing $a=1.32$ fm in the spatial direction. This corresponds to the magic momentum cutoff of $p_{\rm max}\simeq 465\,$MeV that best displays the hidden spin-isospin symmetry of QCD [31].
@@ 3 Results and discussion
As noted before, we employ the same lattice action as in Ref. [18], with the key difference being the non-perturbative treatment of the interactions. Our calculations start with the determination of the ground state energies of 3H and 3He. Using the NLEFT framework, we obtain infinite-volume extrapolated ground state energies of $8.33(2)$ MeV and $7.62(2)$ MeV, respectively, as shown in the left panel of Fig. 2. These values are slightly below the empirical values of $8.48$ MeV and $7.72$ MeV, respectively.
Figure 2: The ground state energies of 3He and 3H nuclei as a function of the box size $L$ (left panel). The trajectory of $f_{D}$$f_{E}$ that leaves these energies invariant (right panel). The dark (light) red/blue bands refer to the 90 (99) % confidence level of the extrapolation. In the right panel the red squares display the result of our calculation and the blue line is a 3rd order spline interpolation to guide the eye.
@@
where the explicit forms of $V_{c_{D}}^{(n)}$ and $V_{c_{E}}^{(n)}$ are given in A. It is important to note that all these potential terms, with their various forms of smearing, are associated with the LECs $c_{D}^{(n)}$ and $c_{E}^{(n)}$, also listed in A, making a direct comparison with the continuum values of these LECs impossible. Here, in Eqs. (6) and (7), we introduce the factors $f_{D}$ and $f_{E}$ to allow us to vary the overall contributions from the one-pion-exchange potential $V_{c_{D}}$ and the SU(4) symmetric short-range potential $V_{c_{E}}$. We then explore the trajectory of $f_{D}$ versus $f_{E}$ that maintains the binding energies of the three-body systems invariant [9, 10]. In the following analysis, we vary $f_{E}$ within the range $[0.4,1.4]$, which, due to the aforementioned correlation, causes $f_{D}$ to vary from $0.7$ to $1.0$, as shown in the right panel of Fig. 2. Because the potential terms in Eqs. (6) and (7) respond differently to the three-body systems, the trajectory observed in Fig. 2 is more complex than the one presented in Ref. [10].
Since no established infinite-volume extrapolation methods exist for matrix elements involving different initial and final states, we employ the well-known plateau method from lattice QCD to extract the values of $\langle F\rangle$ and $\langle GT\rangle$. This approach is illustrated in Fig. 3. Thus, the Fermi operator as well as the Gamow-Teller operator are simultaneously obtained within the consistent framework of NLEFT, and we obtain,
Figure 3: Plots of Fermi (left panel) and Gamow-Teller (right panel) matrix elements at a function of $L$ and the extraction of the corresponding plateaus.
$\langle F\rangle$ $=$ $0.99949(11)\,,$ (8)
$\langle GT\rangle$ $=$ $1.6743(58)\,.$ (9)
Furthermore, in Fig. 4, these matrix elements are shown as a function of $f_{E}$.
@@
Figure 4: Plots of Fermi (left panel) and Gamow-Teller (middle panel) matrix elements and the triton lifetime (right panel) as a function of the parameter $f_{E}$ as defined in Eq. (7). The blue bands represents uncertainties estimated by a time-honoured plateau. The grey band in the middle panel refers to using the averaged empirical values together with the calculated value of $\langle F\rangle$ (left panel). The impact of the uncertainties of $\langle F\rangle$ are marginal as indicated by the black band inside the gray band.
@@
The uncertainty estimate is based on the variation of the operators over the range of $f_{E}$ and $f_{D}$, as shown in Fig. 4, encompassing sets of three-body forces that describe the three-body systems equally well together with the uncertainties induced by the widths of the plateaus in Fig. 3. A similar dependence on the three-body forces was observed in Ref. [10]. Notably, our value of $\langle F\rangle$ is close to the one given in Eq. (3). The corresponding lifetime is given by
$(1+\delta_{R})\,t_{1/2}\,f_{V}=1105.1(74)~{}\text{s},$ (10)
@@
where the uncertainties are also inherited from $K/G^{2}_{V}$, but dominated by the uncertainties for $\langle GT\rangle$. In the right panel of Fig. 4, we show the dependence of the lifetime on the coupling $f_{E}$. Additionally, we calculate $\langle GT_{F}\rangle$ using $\langle F\rangle$ from Eq. (8) and the averaged empirical half-lives from Refs. [23, 24], injected into Eq. (2), giving (see also the middle panel in Fig. 4),
@@ Appendix A Hamiltonian at N3LO
with $\Lambda_{\pi}=300$ MeV the regulator parameter and $M_{\pi}=134.98$ MeV the pion mass. Also, $V_{\rm Coulomb}$ is the Coulomb interaction, $V_{\rm 3N}^{\rm Q^{3}}$ is the 3N potential, $V_{\rm 2N}^{\rm Q^{4}}$ is the 2N short-range interaction at N3LO, $W_{\rm 2N}^{\rm Q^{4}}$ is the 2N Galilean invariance restoration (GIR) interaction at N3LO. For the details of the Coulomb interaction and the two-nucleon (2N) short-range interactions we refer the reader to Ref. [33]. The three-nucleon (3N) interactions at ${\rm Q^{3}}$ consist of locally smeared contact interactions, one-pion exchange interaction with that the two-nucleon contact terms are smeared locally, two-pion exchange potential [7, 9, 36], and two additional SU(4) symmetric potentials denoted by $V_{c_{E}}^{(l)}$ and $V_{c_{E}}^{(t)}$. Therefore, the three-nucleon interactions at ${\rm Q^{3}}$ has the form
@@
Finally, we note that the pion-nucleon vertices are not smeared, so that we can take the values of the dimension-two LECs $4c_{i}$ from Ref. [8].